OpenStudy (anonymous):

Find sin 2x, cos 2x, and tan 2x if tan x= 2/3

OpenStudy (anonymous):

n is in quadrant 1

OpenStudy (jennychan12):

draw a triangle. remember than sin2x = 2sinxcosx cos2x = 1-2sin^2x tan2x = \[\frac{ 2tanx }{ 1-\tan^2x }\]